News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Saturday 11 September 1999 10.00am EST

During he last 24 hours the on-board load continued to operate as expected with a series of HRC off-axis measurements. The load was interrupted at approx 3pm 9/10 in order to allow for a series of diagnostic tests during communication passes to investigate the degradation on the energy resolution of ACIS. The remaining HRC measurements from the load will be rescheduled in a future load.

The ACIS engineering team developed a series of tests including warm boot of ACIS, dumping on-board micro-code, raising the focal plane temperature from -100C to -90C and performing a dark current measurement, power cycling ACIS and adjusting the chip voltages. All tests are expected to be complete prior to entering the radiation zone at approx 1pm 9/11. Analysis of data from the first three tests in is process.

The next mission load (P13) was completed and products distributed for review, this load begins at approx midnight 9/11 and starts with a coordinated observation with HST of Capella.





Return to Status Reports